# On Poincaré–Sobolev level involving fractional GJMS operators on hyperbolic space

### v1

@@ 4.2 Pseudodifferential Operators and Energy Asymptotic
$J_{+}(r)=J_{+}^{(0)}(r)+J_{+}^{(1)}(r),\qquad J_{+}^{(0)}(r):=\int_{\mathbb{R}}e^{i\beta r}\,\widetilde{m}_{s}(\beta)\,\overline{c(\beta)}^{-1}\,d\beta.$
@@
$J_{+}^{(0)}(r)=\int\chi(\beta)a(\beta)e^{i\beta r}\,d\beta+\int(1-\chi(\beta))a(\beta)e^{i\beta r}\,d\beta=:J_{+,\mathrm{low}}^{(0)}(r)+J_{+,\mathrm{high}}^{(0)}(r).$
@@
$J_{+,\mathrm{high}}^{(0)}(r)=\frac{1}{(ir)^{N}}\int\partial_{\beta}^{N}\!\big((1-\chi(\beta))a(\beta)\big)\,e^{i\beta r}\,d\beta,$
@@
For each fixed $r\geq 1$, the series converges uniformly in $\beta$ on compact sets [28, Remark 1 after Theorem 3.1]. Moreover, the coefficients $\Gamma_{k}(\beta)$ have at most polynomial growth in $\beta$ (see the discussion following Theorem 3.1 and the estimates on $\Gamma_{k}$ in Section 3 of [28]), hence the factor $e^{-2r}$ implies that $J_{+}^{(1)}(r)$ enjoys the same oscillatory integration–by–parts bound as $J_{+}^{(0)}(r)$, up to an additional $e^{-2r}$ factor. Concretely, repeating the above argument with the amplitude $a(\beta)\big(\phi_{\beta}(r)-1\big)$ yields
@@
Proof: (i) Fix $\lambda>\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$. By the definition of $\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, there exists $\phi\in C_{c}^{\infty}(\mathbb{H}^{n})\setminus\{0\}$ such that
$Q_{\lambda}(\phi):=\int_{\mathbb{H}^{n}}(\widetilde{\mathcal{P}}_{s}\phi)\,\phi\,dV_{\mathbb{H}^{n}}-\lambda\int_{\mathbb{H}^{n}}|\phi|^{2}\,dV_{\mathbb{H}^{n}}<0.$
Set $q:=-Q_{\lambda}(\phi)>0$ and $p:=2_{s}^{*}$.
Step 1: many far-apart copies. Let $K:=\operatorname{supp}\phi$, which is compact. For each $N\in\mathbb{N}$ we choose isometries $\{\tau_{j}\}_{j=1}^{N}\subset\mathrm{Isom}(\mathbb{H}^{n})$ such that the sets
@@
where $R_{N}>0$ will be chosen later. Define $\phi_{j}:=\phi\circ\tau_{j}^{-1}$ and
$u_{N}:=\sum_{j=1}^{N}\phi_{j}\in C_{c}^{\infty}(\mathbb{H}^{n}).$
@@
$\int_{\mathbb{H}^{n}}|u_{N}|^{2}\,dV_{\mathbb{H}^{n}}=\sum_{j=1}^{N}\int_{\mathbb{H}^{n}}|\phi_{j}|^{2}\,dV_{\mathbb{H}^{n}}=N\int_{\mathbb{H}^{n}}|\phi|^{2}\,dV_{\mathbb{H}^{n}},$ (4.40)
@@
$\int_{\mathbb{H}^{n}}|u_{N}|^{p}\,dV_{\mathbb{H}^{n}}=\sum_{j=1}^{N}\int_{\mathbb{H}^{n}}|\phi_{j}|^{p}\,dV_{\mathbb{H}^{n}}=N\int_{\mathbb{H}^{n}}|\phi|^{p}\,dV_{\mathbb{H}^{n}}.$ (4.41)
@@
$\Bigl(\int_{\mathbb{H}^{n}}|u_{N}|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/p}=N^{2/p}\Bigl(\int_{\mathbb{H}^{n}}|\phi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/p}=N^{\frac{n-2s}{n}}\Bigl(\int_{\mathbb{H}^{n}}|\phi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/p}.$ (4.42)
Step 2: control of the quadratic form. Write $u_{N}=\sum_{j=1}^{N}\phi_{j}$. Expanding the quadratic form gives
$Q_{\lambda}(u_{N})$ $=\sum_{j=1}^{N}Q_{\lambda}(\phi_{j})\;+\;2\sum_{1\leq i<j\leq N}\int_{\mathbb{H}^{n}}(\widetilde{\mathcal{P}}_{s}\phi_{i})\,\phi_{j}\,dV_{\mathbb{H}^{n}}.$
By Lemma 4.4 and $dV_{\mathbb{H}^{n}}$, $Q_{\lambda}(\phi_{j})=Q_{\lambda}(\phi)=-q$, thus $\sum_{j=1}^{N}Q_{\lambda}(\phi_{j})=-Nq$.
By Proposition 4.10, there exist constants $\alpha>0$ and $C>0$ (depending on $n,s$ and $\phi$) such that whenever $\operatorname{dist}(\operatorname{supp}f,\operatorname{supp}g)\geq R$,
@@
Applying (4.43) with $f=\phi_{i}$, $g=\phi_{j}$ and $R=R_{N}$ yields
$\biggl|\sum_{1\leq i<j\leq N}\int_{\mathbb{H}^{n}}(\mathcal{P}_{s}\phi_{i})\,\phi_{j}\,dV_{\mathbb{H}^{n}}\biggr|\leq C\binom{N}{2}e^{-\alpha R_{N}}\leq CN^{2}e^{-\alpha R_{N}}.$
@@
$\frac{Q_{\lambda}(u_{N})}{\bigl(\int_{\mathbb{H}^{n}}|u_{N}|^{p}\,dV_{\mathbb{H}^{n}}\bigr)^{2/p}}\leq-\frac{q}{2}\,\frac{N}{N^{2/p}}\,\Bigl(\int_{\mathbb{H}^{n}}|\phi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{-2/p}\\ =-kN^{1-\frac{2}{p}}=-kN^{\frac{2s}{n}}\longrightarrow-\infty,$
where $k:=\dfrac{q}{2}\Bigl(\int_{\mathbb{H}^{n}}|\phi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{-2/p}>0$. Taking the infimum over $u$ gives $H_{n,s}(\lambda)=$-$\infty$.
@@
(iii) By [24, Theorem 1.9], we obtain $\widetilde{H}_{n,s}\!\bigl(\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\bigr)=S_{n,s}.$ By Proposition 5.3, $\widetilde{H}_{n,s}\!\bigl(0\bigr)=S_{n,s},$ thus we complete the proof.
@@
(iv) By [24, Theorem 1.4] and Proposition 5.3, for $\lambda\leq{\lambda}_{0,s}^{\mathrm{conf}}$, we obtain $H_{n,s}\!\bigl({\lambda}_{0,s}^{\mathrm{conf}}\bigr)=S_{n,s}=H_{n,s}(\lambda).$ The remaining conclusions follow from Proposition 4.11.
@@ 5.1 Stability and Attainability of Sobolev Levels
$\int_{\mathbb{R}^{n}}|\nabla u_{\lambda}|^{2}\,dx=\int_{\mathbb{R}^{n}}|\nabla u|^{2}\,dx,\quad\int_{\mathbb{R}^{n}}|u_{\lambda}|^{2^{*}}\,dx=\int_{\mathbb{R}^{n}}|u|^{2^{*}}\,dx$
@@
Proposition 5.3.
Let $n\geq 2$ and $s\in(0,\tfrac{n}{2})$. Then for all $\lambda\leq 0$,
$H_{n,s}(0)\leq H_{n,s}(\lambda)\leq S_{n,s};\quad\widetilde{H}_{n,s}(\lambda)=\widetilde{H}_{n,s}(0)=S_{n,s}$
where $S_{n,s},H_{n,s}(\lambda),\widetilde{H}_{n,s}(\lambda)$ are defined in (1.12) (1.1)(1.7).
Proof: It is sufficient to show that $H_{n,s}(\lambda)\leq S_{n,s}$ and $\widetilde{H}_{n,s}(\lambda)\leq S_{n,s}$. Fix $p\in\mathbb{H}^{n}$ and choose $0\not\equiv\phi\in C_{c}^{\infty}(B_{1}(0))$. For $\varepsilon\in(0,1)$, in geodesic normal coordinates at $p$, define
$u_{\varepsilon}(x):=\varepsilon^{-\frac{n-2s}{2}}\,\phi\Bigl(\frac{\exp_{p}^{-1}(x)}{\varepsilon}\Bigr),\qquad x\in\mathbb{H}^{n}.$
Write $x=\exp_{p}(\varepsilon z)$, so $y=\exp_{p}^{-1}(x)=\varepsilon z$ and $u_{\varepsilon}(\exp_{p}(\varepsilon z))=\varepsilon^{-\frac{n-2s}{2}}\phi(z)$. In normal coordinates,
hence under $y=\varepsilon z$,
uniformly on the support of $\phi$.
[…]
$\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}|\phi(z)|^{2_{s}^{*}}\,\bigl(1+O(\varepsilon^{2}|z|^{2})\bigr)\,dz=\int_{\mathbb{R}^{n}}|\phi|^{2_{s}^{*}}\,dz+O(\varepsilon^{2}).$
Similarly,
$\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}\varepsilon^{-(n-2s)}|\phi(z)|^{2}\,\bigl(1+O(\varepsilon^{2}|z|^{2})\bigr)\,\varepsilon^{n}\,dz=\varepsilon^{2s}\int_{\mathbb{R}^{n}}|\phi|^{2}\,dz+o(\varepsilon^{2s}).$
Moreover, using the decomposition (1.6), we write
$\Bigl|\int_{\mathbb{H}^{n}}u_{\varepsilon}\,B_{s}(\mathcal{A})\,u_{\varepsilon}\,dV_{\mathbb{H}^{n}}\Bigr|\;\lesssim\;\|u_{\varepsilon}\|_{L^{2}(\mathbb{H}^{n})}^{2}\;=\;\varepsilon^{2s}\int_{\mathbb{R}^{n}}|\phi|^{2}\,dz+o(\varepsilon^{2s}).$
we set
$\int_{\mathbb{H}^{n}}u_{\varepsilon}\,\widetilde{\mathcal{P}}_{s}u_{\varepsilon}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}v_{\varepsilon}\,(-\Delta)^{s}v_{\varepsilon}\,dx=\int_{\mathbb{R}^{n}}\phi\,(-\Delta)^{s}\phi\,dy\;+\;O(\varepsilon^{2}).$
Combining the two pieces yields
$\int_{\mathbb{H}^{n}}u_{\varepsilon}\,\mathcal{P}_{s}u_{\varepsilon}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}\phi\,(-\Delta)^{s}\phi\,dy\;+\;O(\varepsilon^{2})\;+\;O(\varepsilon^{2s})=\int_{\mathbb{R}^{n}}\phi\,(-\Delta)^{s}\phi\,dy\;+\;O\!\bigl(\varepsilon^{2\min\{1,s\}}\bigr).$
Thus, for $\lambda\leq 0$, we have
$\frac{\displaystyle\int_{\mathbb{H}^{n}}(\mathcal{P}_{s}u_{\varepsilon})u_{\varepsilon}\,dV_{\mathbb{H}^{n}}-\lambda\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2}\,dV_{\mathbb{H}^{n}}}{\Bigl(\displaystyle\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/2_{s}^{*}}}$ $=\frac{\displaystyle\Bigl[\int_{\mathbb{R}^{n}}\phi(-\Delta)^{s}\phi\,dy\Bigr]+O(\varepsilon^{2\min\{1,s\}})\;+\;(-\lambda)\,\varepsilon^{2s}\!\int_{\mathbb{R}^{n}}|\phi|^{2}\,dy+o(\varepsilon^{2s})}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|\phi|^{2_{s}^{*}}\,dy+O(\varepsilon^{2})\Bigr)^{\!2/2_{s}^{*}}}$
$\leq\frac{\displaystyle\int_{\mathbb{R}^{n}}\phi(-\Delta)^{s}\phi\,dy}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|\phi|^{2_{s}^{*}}\,dy\Bigr)^{\!2/2_{s}^{*}}}\;+\;C\Bigl(\varepsilon^{2\min\{1,s\}}+|\lambda|\,\varepsilon^{2s}+\varepsilon^{2}\Bigr),$
where $C>0$ depends only on $n,s$ and $\phi$. Letting $\varepsilon\downarrow 0$ and then infimizing over $\phi\in C_{c}^{\infty}(\mathbb{R}^{n})\setminus\{0\}$ gives $H_{n,s}(\lambda)\leq S_{n,s}$. Similarly, we obtain $\widetilde{H}_{n,s}(\lambda)\leq S_{n,s}$. By Proposition 2.3, we have
$\widetilde{H}_{n,s}(\lambda)=\widetilde{H}_{n,s}(0)=S_{n,s}.$
Since $\lambda_{0,s}^{\mathrm{conf}}$ and $\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$ are the bottom of the spectrum of $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s},$ respectively, we immediately obtain the following two results.
@@
Part 2: $s\in(0,1)$:
(i) Using [3, Claim 14.1], Proposition 5.2, Lemma 5.6, the proof is completely analogous to the case $s=1$, and we therefore omit the details.
(ii) By [3, Proposition 16.4], we obtain $(\lambda_{s}^{*},\infty)\subset\mathcal{G}_{n,s}[S_{n,s}]$ and thus, $S_{n,s}(\lambda)$ is attained if $\lambda\in(\lambda_{s}^{*},\infty)$. By Proposition 5.2, Lemma 5.6, for any $\lambda\leq 0<\lambda_{s}^{*}<\mu_{1}<\mu_{2}$,
@@
By the proof of [25, Theorem 1.5, Theorem 1.6], we know that $H_{n,1}(\lambda)$ is achieved if and only if $\lambda\in(0,\lambda_{0,1}^{\mathrm{conf}}].$ Thus, by Lemma 5.6(ii), for any $\lambda\leq 0<\mu_{1}<\mu_{2}<\frac{1}{4},$ we have
@@
(ii) For dimension $n=3$, by [1, Theorem 1.1], we have $H_{3,1}(\frac{1}{4})=S_{3,1}.$ Thus, by Proposition 5.3, we obtain for any $\mu_{1}\leq\frac{1}{4},$

### v2

@@ 4.2 Pseudodifferential Operators and Energy Asymptotic
$J_{+}(r)=J_{+}^{(0)}(r)+J_{+}^{(1)}(r),\qquad J_{+}^{(0)}(r):=\int_{\mathbb{R}}e^{{\bf i}\beta r}\,\widetilde{m}_{s}(\beta)\,\overline{c(\beta)}^{-1}\,d\beta.$
@@
$J_{+}^{(0)}(r)=\int\chi(\beta)a(\beta)e^{{\bf i}\beta r}\,d\beta+\int(1-\chi(\beta))a(\beta)e^{{\bf i}\beta r}\,d\beta=:J_{+,\mathrm{low}}^{(0)}(r)+J_{+,\mathrm{high}}^{(0)}(r).$
@@
$J_{+,\mathrm{high}}^{(0)}(r)=\frac{1}{({\bf i}r)^{N}}\int\partial_{\beta}^{N}\!\big((1-\chi(\beta))a(\beta)\big)\,e^{{\bf i}\beta r}\,d\beta,$
@@
For each fixed $r\geq 1$, the series converges uniformly in $\beta$ on compact sets [40, Remark 1 after Theorem 3.1]. Moreover, the coefficients $\Gamma_{k}(\beta)$ have at most polynomial growth in $\beta$ (see the discussion following Theorem 3.1 and the estimates on $\Gamma_{k}$ in Section 3 of [40]), hence the factor $e^{-2r}$ implies that $J_{+}^{(1)}(r)$ enjoys the same oscillatory integration–by–parts bound as $J_{+}^{(0)}(r)$, up to an additional $e^{-2r}$ factor. Concretely, repeating the above argument with the amplitude $a(\beta)\big(\phi_{\beta}(r)-1\big)$ yields
@@
As a direct consequence of the off-diagonal decay estimate, we obtain the threshold blow-down phenomenon: once $\lambda$ exceeds the spectral bottom, the corresponding Poincaré–Sobolev level collapses to -$\infty$.
@@
Proof: (i) Fix $\lambda>\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$. By the definition of $\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, there exists $\varphi\in C_{c}^{\infty}(\mathbb{H}^{n})\setminus\{0\}$ such that
$Q_{\lambda}(\varphi):=\int_{\mathbb{H}^{n}}(\widetilde{\mathcal{P}}_{s}\varphi)\,\varphi\,dV_{\mathbb{H}^{n}}-\lambda\int_{\mathbb{H}^{n}}|\varphi|^{2}\,dV_{\mathbb{H}^{n}}<0.$
Set $q:=-Q_{\lambda}(\varphi)>0$ and $p:=2_{s}^{*}$.
Step 1: many far-apart copies. Let $K:=\operatorname{supp}\varphi$, which is compact. For each $N\in\mathbb{N}$ we choose isometries $\{\tau_{j}\}_{j=1}^{N}\subset\mathrm{Isom}(\mathbb{H}^{n})$ such that the sets
@@
where $R_{N}>0$ will be chosen later. Define $\varphi_{j}:=\varphi\circ\tau_{j}^{-1}$ and
$u_{N}:=\sum_{j=1}^{N}\varphi_{j}\in C_{c}^{\infty}(\mathbb{H}^{n}).$
@@
$\int_{\mathbb{H}^{n}}|u_{N}|^{2}\,dV_{\mathbb{H}^{n}}=\sum_{j=1}^{N}\int_{\mathbb{H}^{n}}|\varphi_{j}|^{2}\,dV_{\mathbb{H}^{n}}=N\int_{\mathbb{H}^{n}}|\varphi|^{2}\,dV_{\mathbb{H}^{n}},$ (4.40)
@@
$\int_{\mathbb{H}^{n}}|u_{N}|^{p}\,dV_{\mathbb{H}^{n}}=\sum_{j=1}^{N}\int_{\mathbb{H}^{n}}|\varphi_{j}|^{p}\,dV_{\mathbb{H}^{n}}=N\int_{\mathbb{H}^{n}}|\varphi|^{p}\,dV_{\mathbb{H}^{n}}.$ (4.41)
@@
$\Bigl(\int_{\mathbb{H}^{n}}|u_{N}|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/p}=N^{2/p}\Bigl(\int_{\mathbb{H}^{n}}|\varphi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/p}=N^{\frac{n-2s}{n}}\Bigl(\int_{\mathbb{H}^{n}}|\varphi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/p}.$ (4.42)
Step 2: estimates of the quadratic form. Write $u_{N}=\sum_{j=1}^{N}\varphi_{j}$. Expanding the quadratic form gives
$Q_{\lambda}(u_{N})$ $=\sum_{j=1}^{N}Q_{\lambda}(\varphi_{j})\;+\;2\sum_{1\leq i<j\leq N}\int_{\mathbb{H}^{n}}(\widetilde{\mathcal{P}}_{s}\varphi_{i})\,\varphi_{j}\,dV_{\mathbb{H}^{n}}.$
By Lemma 4.4 and $dV_{\mathbb{H}^{n}}$, $Q_{\lambda}(\varphi_{j})=Q_{\lambda}(\varphi)=-q$, thus $\sum_{j=1}^{N}Q_{\lambda}(\varphi_{j})=-Nq$.
By Proposition 4.10, there exist constants $\alpha>0$ and $C>0$ (depending on $n,s$ and $\varphi$) such that whenever $\operatorname{dist}(\operatorname{supp}f,\operatorname{supp}g)\geq R$,
@@
Applying (4.43) with $f=\varphi_{i}$, $g=\varphi_{j}$ and $R=R_{N}$ yields
$\biggl|\sum_{1\leq i<j\leq N}\int_{\mathbb{H}^{n}}(\mathcal{P}_{s}\varphi_{i})\,\varphi_{j}\,dV_{\mathbb{H}^{n}}\biggr|\leq C\binom{N}{2}e^{-\alpha R_{N}}\leq CN^{2}e^{-\alpha R_{N}}.$
@@
$\frac{Q_{\lambda}(u_{N})}{\bigl(\int_{\mathbb{H}^{n}}|u_{N}|^{p}\,dV_{\mathbb{H}^{n}}\bigr)^{2/p}}\leq-\frac{q}{2}\,\frac{N}{N^{2/p}}\,\Bigl(\int_{\mathbb{H}^{n}}|\varphi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{-2/p}\\ =-kN^{1-\frac{2}{p}}=-kN^{\frac{2s}{n}}\to-\infty,$
where $k:=\dfrac{q}{2}\Bigl(\int_{\mathbb{H}^{n}}|\varphi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{-2/p}>0$. Taking the infimum over $u$ gives $H_{n,s}(\lambda)=$-$\infty$.
@@
We next record the basic comparison properties of the hyperbolic fractional Poincaré–Sobolev levels with respect to the Euclidean sharp constant $S_{n,s}$.
Proposition 4.12.
Let $n\geq 2$ and $s\in(0,\tfrac{n}{2})$. Then for all $\lambda\in\mathbb{R}$,
$H_{n,s}(\lambda)\leq S_{n,s};\quad\widetilde{H}_{n,s}(\lambda)\leq S_{n,s}$
where $S_{n,s},H_{n,s}(\lambda),\widetilde{H}_{n,s}(\lambda)$ are defined in (1.12) (1.1)(1.7). In particular, for $\lambda\leq 0,$ we have
$\widetilde{H}_{n,s}(\lambda)=\widetilde{H}_{n,s}(0)=S_{n,s}.$
Proof: It is sufficient to show that $H_{n,s}(\lambda)\leq S_{n,s}$ and $\widetilde{H}_{n,s}(\lambda)\leq S_{n,s}$. Fix $p\in\mathbb{H}^{n}$ and choose $0\not\equiv\varphi\in C_{c}^{\infty}(B_{1}(0))$. For $\varepsilon\in(0,1)$, in geodesic normal coordinates at $p$, define
$u_{\varepsilon}(x):=\varepsilon^{-\frac{n-2s}{2}}\,\varphi\Bigl(\frac{\exp_{p}^{-1}(x)}{\varepsilon}\Bigr),\qquad x\in\mathbb{H}^{n}.$
Write $x=\exp_{p}(\varepsilon z)$, so $y=\exp_{p}^{-1}(x)=\varepsilon z$ and $u_{\varepsilon}(\exp_{p}(\varepsilon z))=\varepsilon^{-\frac{n-2s}{2}}\varphi(z)$. In normal coordinates,
hence under $y=\varepsilon z$,
uniformly on the support of $\varphi$.
[…]
$\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}|\varphi(z)|^{2_{s}^{*}}\,\bigl(1+O(\varepsilon^{2}|z|^{2})\bigr)\,dz=\int_{\mathbb{R}^{n}}|\varphi|^{2_{s}^{*}}\,dz+O(\varepsilon^{2}).$
Similarly,
$\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}\varepsilon^{-(n-2s)}|\varphi(z)|^{2}\,\bigl(1+O(\varepsilon^{2}|z|^{2})\bigr)\,\varepsilon^{n}\,dz=\varepsilon^{2s}\int_{\mathbb{R}^{n}}|\varphi|^{2}\,dz+o(\varepsilon^{2s}).$
Moreover, using the decomposition (1.6), we write
$\Bigl|\int_{\mathbb{H}^{n}}u_{\varepsilon}\,B_{s}(\mathcal{A})\,u_{\varepsilon}\,dV_{\mathbb{H}^{n}}\Bigr|\;\lesssim\;\|u_{\varepsilon}\|_{L^{2}(\mathbb{H}^{n})}^{2}\;=\;\varepsilon^{2s}\int_{\mathbb{R}^{n}}|\varphi|^{2}\,dz+o(\varepsilon^{2s}).$
we set
$\int_{\mathbb{H}^{n}}u_{\varepsilon}\,\widetilde{\mathcal{P}}_{s}u_{\varepsilon}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}v_{\varepsilon}\,(-\Delta)^{s}v_{\varepsilon}\,dx=\int_{\mathbb{R}^{n}}\varphi\,(-\Delta)^{s}\varphi\,dy\;+\;O(\varepsilon^{2}).$
Combining the two pieces yields
$\int_{\mathbb{H}^{n}}u_{\varepsilon}\,\mathcal{P}_{s}u_{\varepsilon}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}\varphi\,(-\Delta)^{s}\varphi\,dy\;+\;O(\varepsilon^{2})\;+\;O(\varepsilon^{2s})=\int_{\mathbb{R}^{n}}\varphi\,(-\Delta)^{s}\varphi\,dy\;+\;O\!\bigl(\varepsilon^{2\min\{1,s\}}\bigr).$
Thus, we have
$\frac{\displaystyle\int_{\mathbb{H}^{n}}(\mathcal{P}_{s}u_{\varepsilon})u_{\varepsilon}\,dV_{\mathbb{H}^{n}}-\lambda\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2}\,dV_{\mathbb{H}^{n}}}{\Bigl(\displaystyle\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/2_{s}^{*}}}$ $=\frac{\displaystyle\Bigl[\int_{\mathbb{R}^{n}}\varphi(-\Delta)^{s}\varphi\,dy\Bigr]+O(\varepsilon^{2\min\{1,s\}})\;+\;(-\lambda)\,\varepsilon^{2s}\!\int_{\mathbb{R}^{n}}|\varphi|^{2}\,dy+o(\varepsilon^{2s})}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|\varphi|^{2_{s}^{*}}\,dy+O(\varepsilon^{2})\Bigr)^{\!2/2_{s}^{*}}}$
$\leq\frac{\displaystyle\int_{\mathbb{R}^{n}}\varphi(-\Delta)^{s}\varphi\,dy}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|\varphi|^{2_{s}^{*}}\,dy\Bigr)^{\!2/2_{s}^{*}}}\;+\;C\Bigl(\varepsilon^{2\min\{1,s\}}+|\lambda|\,\varepsilon^{2s}+\varepsilon^{2}\Bigr),$
where $C>0$ depends only on $n,s$ and $\varphi$. Letting $\varepsilon\downarrow 0$ and then infimizing over $\varphi\in C_{c}^{\infty}(\mathbb{R}^{n})\setminus\{0\}$ gives $H_{n,s}(\lambda)\leq S_{n,s}$. Similarly, we obtain $\widetilde{H}_{n,s}(\lambda)\leq S_{n,s}$. By Proposition 2.3, we have
$\widetilde{H}_{n,s}(\lambda)=\widetilde{H}_{n,s}(0)=S_{n,s},\quad\lambda\leq 0.$
The proof is complete. ∎
Finally, we prove Theorem 1.5 and Theorem 1.6 by combining the strict attainment mechanism established above with the threshold behavior at the spectral bottom.
@@
(iii) By [29, Theorem 1.9], we obtain $\widetilde{H}_{n,s}\!\bigl(\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\bigr)=S_{n,s}.$ By Proposition 4.12, $\widetilde{H}_{n,s}\!\bigl(0\bigr)=S_{n,s},$ thus we complete the proof. $\Box$
@@
(iv) By [29, Theorem 1.4] and Proposition 4.12, for $\lambda\leq{\lambda}_{0,s}^{\mathrm{conf}}$, we obtain $H_{n,s}\!\bigl({\lambda}_{0,s}^{\mathrm{conf}}\bigr)=S_{n,s}=H_{n,s}(\lambda).$ The remaining conclusions follow from Proposition 4.11. $\Box$
@@ 5  Appendix
In this appendix, we collect several properties of Sobolev levels in Euclidean settings. Some of these results are classical, while others are reproved here from a perspective motivated by our observations. Building on these properties, we provide the proofs of Propositions 1.1 and 1.2.
@@ 5.1 Stability and Attainability of Sobolev Levels
We first show that adding a positive lower-order $L^{p}$ perturbation term does not change the optimal Sobolev level.
@@
$\int_{\mathbb{R}^{n}}|\nabla u_{\lambda}|^{2}\,dx=\int_{\mathbb{R}^{n}}|\nabla u|^{2}\,dx,\qquad\int_{\mathbb{R}^{n}}|u_{\lambda}|^{2^{*}}\,dx=\int_{\mathbb{R}^{n}}|u|^{2^{*}}\,dx$
@@
Analogously, in the fractional setting, a positive lower-order $L^{p}$ perturbation still leaves the sharp Sobolev level unchanged.
@@
Since $\lambda_{0,s}^{\mathrm{conf}}$ and $\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$ are the spectral bottoms of $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$, respectively, the following two results are immediate. Their proofs are straightforward and therefore omitted.
@@
We next present a very useful abstract lemma, which links attainment of Sobolev-type levels to a strict comparison (strict monotonicity) of Sobolev-type levels.
@@
Using the preceding results together with classical results on the Brezis–Nirenberg problem in Euclidean space, we now prove Propositions 1.1 and 1.2.
@@
Part II: $s\in(0,1)$:
(i) Using [33, Claim 14.1], Proposition 5.2, Lemma 5.5, the proof is completely analogous to the case $s=1$, and we therefore omit the details.
(ii) By [33, Proposition 16.4], we obtain $(\lambda_{s}^{*},\infty)\subset\mathcal{G}_{n,s}[S_{n,s}]$ and thus, $S_{n,s}(\lambda)$ is attained if $\lambda\in(\lambda_{s}^{*},\infty)$. By Proposition 5.2, Lemma 5.5, for any $\lambda\leq 0<\lambda_{s}^{*}<\mu_{1}<\mu_{2}$,
@@
By the proof of [30, Theorem 1.5, Theorem 1.6], we know that $H_{n,1}(\lambda)$ is achieved if and only if $\lambda\in(0,\lambda_{0,1}^{\mathrm{conf}}].$ Thus, by Lemma 5.5(ii), for any $\lambda\leq 0<\mu_{1}<\mu_{2}<\frac{1}{4},$ we have
@@
(ii) For dimension $n=3$, by [2, Theorem 1.1], we have $H_{3,1}(\frac{1}{4})=S_{3,1}.$ Thus, by Proposition 4.12, we obtain for any $\mu_{1}\leq\frac{1}{4},$